Suppose that there are only two periods left in history. Let the marginal net benefit function for iron ore be 180 - 3Q, where Q is the quantity of iron ore consumed in a period. If there are 60 units of iron ore left to use, and the discount rate is 50%, what is the net benefit in each period if the iron ore is divided equally between the two periods? 8100 7992 6840 6750 5400 5184 4536 O 4182 4050 3456 2700 2592 0 1944 1350 864
Answer #1

When iron ore is divided equally between two periods then we have Q1 = Q2 = 30 units

In this case there is no marginal cost and there is no use of discount rate when equal quantities are used in each period. Hence the area under the marginal net benefit curve will be total net benefit

Net benefit = 0.5*(max price - minimum price)*quantity each period = 0.5*(180 - 0)*30 = 2700

Select 2700.